OpenStudy (princessaurora):

**WILL FAN & MEDAL** I need help with this problem The test to detect the presence of a liver disorder is 98% accurate for a person who has the disease and 97% accurate for a person who does not have the disease. If 3.5% of the people in a given population actually have the disorder, what is the probability that a randomly chosen person tests positive? A) 0.0343 B) 0.035 C) 0.06325 D) 0.02895

OpenStudy (astrophysics):

Hii!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

lets do it with numbers !

OpenStudy (misty1212):

or not maybe @Astrophysics has a better way

OpenStudy (astrophysics):

It's cool you can go ahead :)

OpenStudy (misty1212):

i would start by saying lets test 10,000 people

OpenStudy (misty1212):

\[3.5\%\] of them have the disease, and \[3.5\%\] of \(10,000\) is \(350\) so \(350\) people have it

OpenStudy (misty1212):

\[98\%\] of those 350 will test positive, and \[0.98\times 350=343\]so we know that 343 people who have it will test positive

OpenStudy (misty1212):

now lets see how many of those who do not have it will test positive first off, 350 have it so \[10,000-350=9650\] so not have the disease of those \(3\%\) will also test positive

OpenStudy (misty1212):

\[.03\times 9650=289.5\] so \[289.5\] will of those who do not have it will also test positive

OpenStudy (misty1212):

all together, how many people test positive?

OpenStudy (misty1212):

i.e. add \[289.5+343\]

OpenStudy (princessaurora):

632.5?

OpenStudy (misty1212):

ok, then divide by \(10,000\) by moving the decimal over to get your answer

OpenStudy (princessaurora):

0.06325?

OpenStudy (misty1212):

oh imagine that! the answer is once again C when in doubt, charlie out

OpenStudy (princessaurora):

Thanks a lot for your help, you explained it very well :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]